OpenStudy (anonymous):

Solve (3a-2)=a+25/a+1

OpenStudy (anonymous):

3a - 2 = a + 25/a + 1 Bring everything on one side 2a - 25/a - 3 = 0 2a^2 - 3a - 25 = 0 (assume a different from 0) The discriminant is d = 9 + 200 = 209 So the two solutions are a = (3 - sqrt(209))/4 and a = (3 + sqrt(209))/4 sqrt = square root Please check the algebra, I have done it in my mind :)